# RadialRouter: Structured Representation for Efficient and Robust Large Language Models Routing

### v1

@@ Abstract
Ruihan Jin   Pengpeng Shao   Zhengqi Wen   Jinyang Wu Mingkuan Feng   Shuai Zhang   Jianhua Tao Department of Automation, Tsinghua University, Beijing, China {jinrh24, wu-jy23, fmk24}@mails.tsinghua.edu.cn    shuai_zhang@mail.tsinghua.edu.cn {ppshao, zqwen, jhtao}@tsinghua.edu.cn
@@ 1 Introduction
In this paper, we propose a novel LLM routing approach named RadialRouter, which leverages a Transformer-based architecture to enhance performance for efficient and robust LLM routing. To represent the interrelationship between the query and LLMs, we propose RadialFormer as the backbone of RadialRouter and incorporate the structure consisting of a relay node and $n$ satellite nodes. During update, these nodes are processed through the multi-head attention mechanism. Compared with the standard Transformer, RadialFormer reduces the computational complexity from $O(n^{2}d)$ to $O(nd)$. The optimal LLM selection is performed based on the final states of satellite nodes. To guide the selection and enhance comprehensive representation, we employ a Kullback-Leibler divergence loss for supervision. Additionally, we cluster the queries into groups and introduce a query-query contrastive loss, which fosters the generation of similar embeddings for semantically related queries, facilitating robust LLM routing.
@@ 3.1 RadialFormer Architecture
The updated satellite nodes are regularized through layer normalization Ba et al. (2016). The update algorithm of RadialFormer is shown in Alg.1. Given the sequence length $n$ and the dimension of the hidden state $d$, RadialFormer reduces the computational complexity of the standard Transformer from $O(n^{2}d)$ to $O(nd)$. The specific design of RadialFormer integrates both lightweight and structured representations of the routing. Subsequent experiments validate that RadialFormer facilitates efficient and effective LLM routing.
@@ 3.2 Optimal LLM Selection
Given a query $\mathrm{x}_{j}$, the process of LLM selection is supervised by the scores of the candidate LLMs $\mathrm{score}_{\mathrm{x}_{j}}=\{s_{j}^{(i)}:i=1,\dots,n\}$, which we identify in advance (described in Sec.4.3). We employ the Kullback-Leibler divergence loss Kullback and Leibler (1951) for supervision to guide the routing probability toward the probability derived from the exponent of true scores, which is defined as
@@
Input: number of layers $T$, model embeddings $\mathbf{m}_{1},\cdots,\mathbf{m}_{n}$, and query embedding $\mathbf{q}$.
@@
2:$\mathbf{s}^{0}_{1},\cdots,\mathbf{s}^{0}_{n}\leftarrow\mathbf{m}_{1},\cdots,\mathbf{m}_{n}$
3:$\mathbf{r}^{0}\leftarrow\mathbf{q}$
@@ 3.3 Optimization with Contrastive Loss
To enhance the robustness of LLM routing, we introduce a query-query contrastive loss, which promotes the ability of the language encoder in RadialRouter to generate analogous embeddings for semantically similar queries. Following Chen et al. (2024), we transform the query embeddings encoder by a pre-trained language model into low-dimensional vectors by the t-SNE algorithm Van der Maaten and Hinton (2008) and perform $k$-means clustering MacQueen (1967) to obtain $N$ semantic groups $\{\mathcal{K}_{1},\cdots,\mathcal{K}_{N}\}$. We use the sample-sample contrastive loss to promote the generation of embeddings, formulated as:
@@
Finally, we learn the RadialRouter by minimizing a final objective that combines the KL divergence and the query-query contrastive loss as
$\bm{\theta}^{*}=\arg\min\mathbb{E}_{\mathrm{x}\sim\mathcal{D}_{\mathrm{train}}}\mathcal{L}_{\text{KL}}(\mathrm{x};\bm{\theta})+\lambda\mathcal{L}_{\text{q-q}}(\mathrm{x};\bm{\theta}),$ (6)
@@ 4.3 Metrics
  • Cost refers to the average LLM inference cost for generating responses to the queries, which is expressed in dollars per 1M tokens. The statistics of candidate LLMs on RouterBench are shown in Appendix C.
@@ 5.2 Ablation Studies
We conduct ablation studies to investigate the contribution of each component in RadialRouter as shown in Tab.2. Here, ‘$w/o$’ denotes variants with specific components removed, and ‘$+$’ denotes replacing RadialFormer with alternative architectures.
Replacing RadialFormer with alternative architectures (Star-Transformer Guo et al. (2019), standard Transformer Vaswani et al. (2017), and MLP) all leads to performance degradation, indicating the effectiveness of the attention mechanism design in RadialFormer for the structured representation. In construct, the architecture of Star-Transformer introduces unnecessary connections (e.g., Ring Connections), causing interference to the routing process. We further compare the routing efficiencies of different architectures. RadialFormer demonstrates lower time consumption than both Star-Transformer and standard Transformer, validating the effectiveness of its lightweight design.
@@
RadialRouter $w/\ \mathcal{L}_{\text{KL}}$ 0.816 0.757 0.715
RadialRouter $w/\ \mathcal{L}_{\text{ce}}$ 0.533 0.530 0.520
RadialRouter $w/\ \mathcal{L}_{\text{q-L}}$ 0.815 0.714 0.676
Table 4: Comparison on routing to an increasing number of candidate LLMs in the Balance scenario. The results are taken as the average of each dataset.
@@ 5.3.1 Loss Function for LLM Selection
Viewing LLM routing as a multi-class classification problem, the cross-entropy loss is introduced. In this approach, the LLM that receives the highest true score is assigned ‘1’, while all other LLMs are assigned ‘0’. The cross-entropy loss function is defined as
@@
Considering the objective of routing is to allocate the query to top-performing LLMs, rather than merely identifying the optimal model, Chen et al. (2024) introduces the sample-LLM contrastive loss to learn the router. We make minor modifications to it under the RadialRouter framework. Specifically, based on the true scores, we construct the LLMs index set $\mathcal{I}^{+}$ and $\mathcal{I}^{-}$ as the indices of LLMs corresponding to the top-$K$ and bottom-$K$ scores, respectively. The query-LLM contrastive loss is then defined as
$\mathcal{L}_{\text{q-L}}(\mathrm{x};\bm{\theta})=\sum_{i^{+}\in\mathcal{I}^{+}}-\log\frac{e^{p_{i^{+}}}}{e^{p_{i^{+}}}+\sum\limits_{i^{-}\in\mathcal{I}^{-}}e^{p_{i^{-}}}},$ (9)
where $p$ denotes the predicted routing probability.
Tab.3 displays the results of comparison. Training RadialRouter with $\mathcal{L}_{\text{KL}}$ yields the highest scores in all three scenarios, as can be observed. From a macroscopic perspective, $\mathcal{L}_{\text{KL}}$, $\mathcal{L}_{\text{ce}}$ and $\mathcal{L}_{\text{q-L}}$ address the routing problem out of distinct viewpoints: probabilistic distribution fitting, multi-class classification, and similarity matching, respectively. The superiority of the KL divergence underscores the effectiveness of framing the routing problem through the lens of probabilistic distribution fitting, which fosters a comprehensive understanding of the intrinsic connection between the query and LLMs in the routing process, as opposed to focusing merely on a limited number of dominant LLMs. This perspective aligns seamlessly with the design principles of RadialFormer, allowing for the accomplishment of both efficient and robust routing.
@@ 5.3.4 Effects of $\lambda$ on Optimization
We study the impact of the contrastive loss weight $\lambda$ in Eq.6 on the optimization of RadialRouter. Experiments are conducted in the Balance scenario, and the results are presented in Appendix F. As can be seen, the highest score is achieved when $\lambda=0.5$. Therefore, we fix $\lambda$ to 0.5 when training RadialRouter. Moreover, we can observe that RadialRouter demonstrates insensitivity across a wide range of $\lambda\in[0.25,5]$, providing more flexibility for the selection of $\lambda$ in practice.
@@ Appendix A Details of Multi-Head Attention
where $[Q;K;V]=[\mathbf{q}W_{q};\mathbf{H}W_{k};\mathbf{H}W_{v}]$, and $W_{q},W_{k},W_{v}$ denote learnable parameters.
@@
$\mathrm{MHAttn}(\mathbf{q},\mathbf{H})=[\mathrm{head}_{1},\cdots,\mathrm{head}_{h}]W_{o},$ (11)
@@
where $[Q_{i};K_{i},;V_{i}]$ are the $i$-th group from $[Q;K;V]$ with a dimension of $d/h$, and $W_{o}$ denotes output learnable parameter.
@@ Appendix B Training and Inference Algorithm of RadialRouter
1:training set $\mathcal{D}_{\text{train}}$, LLMs $\{\mathrm{LLM}_{i}:i=1,\dots,n\}$, number of out-group queries $H$, number of clusters $N$, hyper-parameter $\lambda$, mini-batch size $b$, and learning rate $\eta$; learnable parameters $\bm{\theta}$: encoder $\mathcal{E}$, RadialFormer $\mathcal{RF}$, MLP $\mathcal{M}$, and learnable LLM embeddings $\{\mathbf{m}_{i}:i=1,\dots,n\}$;
@@
5:Cluster training queries $\{\mathrm{x}_{j}:j=1,\dots,l\}$ into $N$ groups $\{\mathcal{K}_{1},\cdots,\mathcal{K}_{N}\}$;
@@
11:         $\mathbf{r}^{T},\mathbf{s}_{1}^{T},\cdots,\mathbf{s}_{n}^{T}\leftarrow\mathcal{RF}(\mathbf{q},\mathbf{m}_{1},\cdots,\mathbf{m}_{n})$;
@@
22:$i^{\prime}\leftarrow\arg\max_{i\in\{1,\dots,n\}}(p_{i}^{\prime})$;
23:$\hat{\mathrm{y}}^{\prime}\leftarrow\mathrm{LLM}_{i^{\prime}}(\mathrm{x}^{\prime})$.
@@ Appendix F Detailed Results of Effects of $\lambda$ on Optimization
Tab.6 shows the effects of the contrastive loss weight $\lambda$ on the optimization of RadialRouter. Experiments are conducted in the Balance scenario, visualized in Fig.6. As can be seen, the highest score is achieved when $\lambda=0.5$ and RadialRouter demonstrates insensitivity across a wide range of $\lambda\in[0.25,5]$, offering greater flexibility for the selection of $\lambda$ in practice.

### v2

@@ Abstract
Ruihan Jin1   Pengpeng Shao1thanks: Corresponding authors.  Zhengqi Wen [^1] footnotemark: 1  Jinyang Wu1 Mingkuan Feng1   Shuai Zhang1   Jianhua Tao1,2 1Department of Automation, Tsinghua University 2Beijing National Research Center for Information Science and Technology jinrh24@mails.tsinghua.edu.cn   {ppshao, zqwen}@tsinghua.edu.cn
@@ 1 Introduction
In this paper, we propose a novel LLM routing approach named RadialRouter, which leverages a Transformer-based architecture to enhance performance for efficient and robust LLM routing. To represent the interrelationship between the query and LLMs, we propose RadialFormer as the backbone of RadialRouter and incorporate the structure consisting of a relay node and $n$ satellite nodes. During update, these nodes are processed through the multi-head attention mechanism. Compared with the standard Transformer, RadialFormer reduces the computational complexity from $O(l^{2}d)$ to $O(ld)$ given the sequence length $l$ and the dimension of the hidden state $d$. The optimal LLM selection is performed based on the final states of satellite nodes. To guide the selection and enhance comprehensive representation, we employ a Kullback-Leibler divergence loss for supervision. Additionally, we cluster the queries into groups and introduce a query-query contrastive loss, which fosters the generation of similar embeddings for semantically related queries, facilitating robust LLM routing.
@@ 3.1 RadialFormer Architecture
The updated satellite nodes are regularized through layer normalization Ba et al. (2016). The update algorithm of RadialFormer is shown in Alg.1. Given the sequence length $l$ and the dimension of the hidden state $d$, RadialFormer reduces the computational complexity of the standard Transformer from $O(l^{2}d)$ to $O(ld)$. The specific design of RadialFormer integrates both lightweight and structured representations of the routing. Subsequent experiments validate that RadialFormer facilitates efficient and effective LLM routing.
@@ 3.2 Optimal LLM Selection
Given a query $\mathrm{x}_{j}$, the process of LLM selection is supervised by the scores of the candidate LLMs $\mathrm{score}_{\mathrm{x}_{j}}=\{s_{j}^{(i)}:i=1,\dots,n\}$, which we identify in advance (described in Sec.4.3). We employ the Kullback-Leibler divergence loss Kullback and Leibler (1951) for supervision to guide the routing probability toward the probability derived from the exponent of true scores, which is defined as:
@@
Input: number of layers $T$, model embeddings $\mathbf{m}_{1},\ldots,\mathbf{m}_{n}$, and query embedding $\mathbf{q}$.
@@
[…]
@@ 3.3 Optimization with Contrastive Loss
To enhance the robustness of LLM routing, we introduce a query-query contrastive loss, which promotes the ability of the language encoder in RadialRouter to generate analogous embeddings for semantically similar queries. Following Chen et al. (2024), we transform the query embeddings encoder by a pre-trained language model into low-dimensional vectors by the t-SNE algorithm Van der Maaten and Hinton (2008) and perform $k$-means clustering MacQueen (1967) to obtain $N$ semantic groups $\{\mathcal{K}_{1},\ldots,\mathcal{K}_{N}\}$. We use the sample-sample contrastive loss to promote the generation of embeddings, formulated as:
@@
Finally, we learn the RadialRouter by minimizing a final objective that combines the KL divergence and the query-query contrastive loss as:
$\bm{\theta}^{*}=\arg\min\underset{{\mathrm{x}\sim\mathcal{D}_{\mathrm{train}}}}{\mathbb{E}}\mathcal{L}_{\text{KL}}(\mathrm{x};\bm{\theta})+\lambda\mathcal{L}_{\text{q-q}}(\mathrm{x};\bm{\theta}),$ (6)
@@ 4.3 Metrics
  • Cost refers to the average LLM inference cost for generating responses to the queries, which is expressed in dollars. The statistics of candidate LLMs on RouterBench are shown in Appendix C.
@@ 5.2 Ablation Studies
We conduct ablation studies to investigate the contribution of each component in RadialRouter as shown in Tab.2. Here, ‘$w/o$’ denotes variants with specific components removed, and ‘++’ denotes replacing RadialFormer with alternative architectures.
We model the routing problem as a probability distribution prediciton problem which requires efficient representation of both queries and LLMs. From Tab.2 we can observe that replacing RadialFormer with alternative architectures (Star-Transformer Guo et al. (2019), standard Transformer Vaswani et al. (2017), and MLP) all leads to performance degradation. The result indicates the effectiveness of RadialFormer, which maintains rich semantic information through structured representation and facilitates the selection of the optimal LLM. In construct, the architecture of Star-Transformer introduces unnecessary connections (e.g., Ring Connections), causing interference to the routing process. We further compare the routing efficiencies of different architectures. RadialFormer demonstrates lower time consumption than both Star-Transformer and standard Transformer, validating its lightweight design.
@@
[…]
Table 4: Comparison on routing to an increasing number of candidate LLMs in the Balance scenario, where ‘+’ denotes an increase in the LLM pool. The results are taken as the average of each dataset.
@@ 5.3.1 Loss Function for LLM Selection
Viewing LLM routing as a multi-class classification problem, the cross-entropy loss is introduced. In this approach, the LLM that receives the highest true score is assigned ‘1’, while all other LLMs are assigned ‘0’. The cross-entropy loss function is defined as:
@@
Considering the objective of routing is to allocate the query to top-performing LLMs, rather than merely identifying the optimal model, Chen et al. (2024) introduces the sample-LLM contrastive loss to learn the router. We make minor modifications to it within the RadialRouter framework. Based on the true scores obtained by Eq.7, we construct the LLMs index set $\mathcal{I}$ and $\bar{\mathcal{I}}$ as the indices of LLMs corresponding to the top-$K$ and bottom-$K$ scores, respectively. The query-LLM contrastive loss is then defined as:
$\mathcal{L}_{\text{q-L}}(\mathrm{x};\bm{\theta})=\sum_{i\in\mathcal{I}}-\log\frac{e^{p_{i}}}{e^{p_{i}}+\sum\limits_{j\in\bar{\mathcal{I}}}e^{p_{j}}},$ (9)
where $p$ denotes the predicted routing probability, $i$ and $j$ denote the index of LLMs corresponding to the top-$K$ and bottom-$K$ scores, respectively.
Tab.3 displays the results of comparison. Training RadialRouter with $\mathcal{L}_{\text{KL}}$ yields the highest scores in all three scenarios, as can be observed. From a macroscopic perspective, $\mathcal{L}_{\text{KL}}$, $\mathcal{L}_{\text{ce}}$ and $\mathcal{L}_{\text{q-L}}$ address the routing problem out of distinct viewpoints: probabilistic distribution fitting, multi-class classification, and contrastive learning, respectively. The superiority of the KL divergence underscores the effectiveness of framing the routing problem through the lens of probabilistic distribution fitting, which fosters a comprehensive understanding of the intrinsic connection between the query and LLMs in the routing process, as opposed to focusing merely on a limited number of dominant LLMs. This perspective aligns seamlessly with the design principles of RadialFormer, allowing for the accomplishment of both efficient and robust routing.
@@ 5.3.4 Effects of $\lambda$ on Optimization
We study the impact of the contrastive loss weight $\lambda$ in Eq.6 on the optimization of RadialRouter. Experiments are conducted in the Balance scenario, and the results are presented in Appendix F. As can be seen, the highest score is achieved when $\lambda=0.5$. Therefore, we fix $\lambda$ to 0.5 when training RadialRouter. Moreover, we observe that RadialRouter demonstrates insensitivity across a broad range of $\lambda\in[0.25,5]$, which confirms the robustness of our method and provides greater flexibility in the practical selection of $\lambda$.
@@ Appendix A Details of Multi-Head Attention
where $Q;K;V]=[\mathbf{q}W_{q};\mathbf{H}W_{k};\mathbf{H}W_{v}]$, and $W_{q},W_{k},W_{v}$ denote learnable parameters.
@@
$\mathrm{MHAttn}(\mathbf{q},\mathbf{H})=[\mathrm{head}_{1},\ldots,\mathrm{head}_{h}]W_{o},$ (11)
@@
where $Q_{i};K_{i},;V_{i}]$ are the $i$-th group from $[Q;K;V]$ with a dimension of $d/h$, and $W_{o}$ denotes output learnable parameter.
@@ Appendix B Training and Inference Algorithm of RadialRouter
1:training set $\mathcal{D}_{\text{train}}$, LLMs $\{\mathrm{LLM}_{i}:i=1,\dots,n\}$, number of out-group queries $H$, number of clusters NN, hyper-parameter $\lambda$, mini-batch size $b$, and learning rate $\eta$; learnable parameters $\bm{\theta}$: encoder $\mathcal{E}$, RadialFormer $\mathcal{RF}$, MLP $\mathcal{M}$, and learnable LLM embeddings $\{\mathbf{m}_{i}:i=1,\dots,n\}$;
@@
5:Cluster training queries $\{\mathrm{x}_{j}:j=1,\dots,l\}$ into $N$ groups $\{\mathcal{K}_{1},\ldots,\mathcal{K}_{N}\}$;
@@
11:   $\mathbf{r}^{T},\mathbf{s}_{1}^{T},\ldots,\mathbf{s}_{n}^{T}\leftarrow\mathcal{RF}(\mathbf{q},\mathbf{m}_{1},\ldots,\mathbf{m}_{n})$;
@@
[…]
@@ Appendix F Detailed Results of Effects of $\lambda$ on Optimization
Tab.6 shows the effects of the contrastive loss weight $\lambda$ on the optimization of RadialRouter. Experiments are conducted in the Balance scenario, visualized in Fig.6. As can be seen, the highest score is achieved when $\lambda=0.5$ and RadialRouter demonstrates insensitivity across a wide range of $\lambda\in[0.25,5]$, which confirms the robustness of our method and provides greater flexibility for the selection of $\lambda$ in practice.